Lisa Simpson wants to have ​$1,900,000 in 60 years by making equal annual​ end-of-the-year deposits into a​ tax-deferred account paying 9.75 percent annually. What must​ Lisa's annual deposit​ be?
Answer #1

Desired sum at retirement = $1,900,000
Time to retirement = 60 years
Annual interest rate = 9.75%

Let annual deposit be $x

$x*1.0975^59 + $x*1.0975^58 + … + $x*1.0975 + $x = $1,900,000
$x * (1.0975^60 - 1) / 0.0975 = $1,900,000
$x * 2,714.12005 = $1,900,000
$x = $700.04

So, annual deposit is $700.04